OpenStudy (anonymous):

Post a 50-word response to the following question: · What is the Pythagorean theorem? How is it used?

OpenStudy (anonymous):

try wikipidea or google

hero (hero):

50-word response? You need 50 words to explain it? I can't figure out if that's too little or too much

OpenStudy (anonymous):

lol it can be more but not less lol

hero (hero):

I say just explain it and don't worry about how many words. Screw the teacher.....Oh wait...

OpenStudy (anonymous):

she just wants a simple explaination lol

OpenStudy (anonymous):

In any right angle triangle, the square of hypotenus is equal to the sum of the quare of the first right side and the square of the second right side of the right angle.

OpenStudy (anonymous):

hope that's enogh

OpenStudy (anonymous):

i think it needs a little more let me double check though! thnx :)

OpenStudy (anonymous):

ya it needs just a little more detail... like 20 words lol any suggestions??

hero (hero):

You can always include an example

hero (hero):

I wouldn't explain it without at least including this: \[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

leg^2 +leg^2=hypotenuse^2

OpenStudy (anonymous):

Idk how 2 use less!

OpenStudy (anonymous):

ok add this, if ABC is a right angle triangle in , then AC^2=AB^2+BC^2

OpenStudy (anonymous):

I'll delete it...

OpenStudy (anonymous):

so wats the final answer ganna look like??

hero (hero):

raheen, I've never seen pythagorean theorem expressed in that form

OpenStudy (saifoo.khan):

Still Stuck?

OpenStudy (anonymous):

lol yes :D

OpenStudy (anonymous):

im not up to 50 words yet lol

OpenStudy (saifoo.khan):

lol, what you wrote?

OpenStudy (anonymous):

In any right angle triangle, the square of hypotenus is equal to the sum of the quare of the first right side and the square of the second right side of the right angle. ex: AC^2=AB^2+BC^2

OpenStudy (saifoo.khan):

Pythagoras therom is used to find any unknown side of a Right Angle Triangle. A right angle triangle is one which consists of one-90* angle among 3. The opposite side to it is known as Hypotenuse. The formula for solving it is: Base^2 + Perpendicular^2 = Hypotenuse^2 We insert the values and get the respected value from it.